OpenStudy (anonymous):

solve for x. Enter your answer in interval notation using grouping symbols. x^2+5x<24

OpenStudy (anonymous):

This is a quadratic inequality because of the x^2 so pretend it an equation, set it equal to 0 and solve be either factoring or quad formula x^2 +5x - 24 = 0 (x+8)(x-3) = 0 so it is = 0 at x = -8 and x = 3 When is the inequality less than 0 (or negative) We can test numbers like -10000 because that is to the left of -8 (-10000 + 8)(-1000 - 3) = a + number so it is above the x axis a number between -8 and 3 is 0, so (0+8)(0 - 3) is a neg number so the graph is below the x axis. and a number bigger ( to the right ) than 3 is like 10000 (1000+8(1000-3) is + so above the x axis so the solution to the inequality is (-8, 3) in interval notation. Is that the "grouping symbols" you need?